OpenStudy (anonymous):

limx→-3 (√2x+22)-4/x+3

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \lim_{x \rightarrow -3} ~~\frac{\sqrt{2x+22}-4}{x+3} }\)

OpenStudy (solomonzelman):

yes ?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

What have you tried to do ? (On this site we need to provide guidance not answers)

OpenStudy (solomonzelman):

It is not the same on both sides, so the limit exists or not ?

OpenStudy (anonymous):

exists? i think

OpenStudy (solomonzelman):

No it does not.. perhaps it is +4 in your question (not -4) ?

OpenStudy (anonymous):

so how does it not exist? how am i supposed to show work for it haha

OpenStudy (anonymous):

\[\lim_{x \rightarrow -3}\frac{\sqrt{2x+22}-4}{x+3}\cdot\frac{\sqrt{2x+22}+4}{\sqrt{2x+22}+4}\\ \lim_{x \rightarrow -3}\frac{2x+22-16}{(x+3)(\sqrt{2x+22}-4)}\\ \lim_{x \rightarrow -3}\frac{2x+6}{(x+3)(\sqrt{2x+22}-4)}\\ \lim_{x \rightarrow -3}\frac{2(x+3)}{(x+3)(\sqrt{2x+22}-4)}\\ \lim_{x \rightarrow -3}\frac{2}{\sqrt{2x+22}-4}\]

OpenStudy (solomonzelman):

But it is not equal for left and right side? 2 sided limit does not exist.

OpenStudy (solomonzelman):

With +4 on the top, instead of -4, it would okay though.

OpenStudy (solomonzelman):

√(2x+22)-4 → √(2{-3}+22)-4 → √(-6+22)-4 → √16 - 4 denominator = 0 or -8

OpenStudy (solomonzelman):

oops :)

OpenStudy (solomonzelman):

I would like to know how though

OpenStudy (anonymous):

I think you're using a slightly incorrect conjugate? We're given - in the numerator, so you multiply by +.

OpenStudy (solomonzelman):

I am not using a conjugate, I am continuing from where you left off (calculations in black latex) and just plugging -3 for x.

OpenStudy (anonymous):

Oh I see the mistake, *I* was using the wrong conjugate :P The last line should be \[\lim_{x\to-3}\frac{2}{\sqrt{2x+22}~{\color{red}{\Huge{+}}}~4}\] Thanks for catching that.

OpenStudy (solomonzelman):

yes, but for √16 = -4 as well as +4, you still get denominator = 0 and 8 (instead of -8)

OpenStudy (solomonzelman):

The highest math course I took was trig, so I don't really get how is it 1/4. Wolfrrrrram

OpenStudy (anonymous):

\[x^2=16~~\Rightarrow~~x=\pm\sqrt{16}=\pm4\] but we're already given the positive square root, \(\sqrt{16}=4\).

OpenStudy (anonymous):

Wolfram says the same: http://www.wolframalpha.com/input/?i=Sqrt%5B16%5D%3D-4

OpenStudy (solomonzelman):

yes, so with the positive root... I agree

OpenStudy (solomonzelman):

Do you always take the positive root, when solving such limits ?

OpenStudy (anonymous):

Well, if the expression contains \(\sqrt{}\) alone (without a sign in front) we assume the positive root. The root is only negative if there's a minus sign in front. Compare: \[\sqrt {16}+1=4+1=5\\ 1-\sqrt {9}=1+(-3)=1-3=-2\]

OpenStudy (solomonzelman):

Yes... ty! I'll definitely remember that, I'll need it in calculus.

OpenStudy (anonymous):

No problem. Good luck

OpenStudy (solomonzelman):

:)